OpenStudy (anonymous):

If the Harmonic mean b/w two +ve Numbers is to their G.M as 12:13 then numbers are in ratio?

OpenStudy (anonymous):

@mukushla @theyatin @ghazi

OpenStudy (anonymous):

h.m = ab/a+b g.m =\[\sqrt{ab}\]

OpenStudy (anonymous):

h.m = 2ab/a+b....

OpenStudy (anonymous):

ohk sorry typo

OpenStudy (anonymous):

then divide h.m by g.m equate it with 12/13 solve to get answer

OpenStudy (anonymous):